# Leveraging Online Data to Enhance Medical Knowledge in a Small Persian Language Model

### v2

@@ VI-B Comparison with pipeline alternatives
As previously mentioned, one alternative to creating a Persian medical language model is a pipeline system. However, a major problem with pipeline systems is their speed; they exhibit high inference times because the output from one model must be fed into a second model, and then the output of the second model is processed again by the first model. This iterative process significantly hampers efficiency. To address the low speed of pipeline models, we have loaded all parameters—both those pertaining to the translators and the medical language model—simultaneously. Our experiments with models such as Medmobile [9] paired with gemma-2b-it [14] as translators, and Medmobile paired with parsinlu [32] [34] models, showed disappointing results, as evidenced by the poor performance displayed in Table IV.
Another significant issue with the pipeline alternative is its poor performance in accurately detecting and translating medical terms. This limitation poses a serious challenge, as precise terminology is crucial for effective communication in healthcare settings, where misunderstandings can have serious consequences for patient care and treatment outcomes. The underlying cause of this deficiency is likely due to the fact that the translators employed in these pipeline systems have not been specifically developed for medical translation. Unlike general-purpose translation models, medical translation requires a nuanced understanding of specialized vocabulary, context, and the intricacies of medical language. Currently, there are no models tailored for medical translation in the Persian language, which means that existing systems are ill-equipped to handle the complexities of medical terminology. As shown in Figure 4, these limitations have resulted in all pipeline alternatives achieving lower win rates against our model, Gaokerena. This performance gap highlights the inadequacies of current pipeline approaches in meeting the demands of medical translation, emphasizing the necessity for the development of dedicated medical translator models that can effectively address these specific challenges.
TABLE IV: our model performance in comparison with pipeline alternatives
Gao- MedMobile MedMobile
kerena + gemma2 + parsinlu
@@
As mentioned in the baseline model section, we can merge our updated parameters into aya-vision instead of aya-exapanse, This will allow us to incorporate medical images, such as MRIs and CT scans, as inputs to our model. However, it is important to note that our models capabilities with multi-modal queries are under-explored. our future research plans involve to exploring and enhancing our model’s responses to queries that are accompanied by medical images.

### v3

@@ VI-B Comparison with pipeline alternatives
As previously mentioned, one alternative to creating a Persian medical language model is a pipeline system. However, a major problem with pipeline systems is their speed; they exhibit high inference times because the output from one model must be fed into a second model, and then the output of the second model is processed again by the first model. This iterative process significantly hampers efficiency. To address the low speed of pipeline models, we have loaded all parameters—both those pertaining to the translators and the medical language model—simultaneously.
Our experiments with models such as Medmobile [9] paired with gemma-2b-it [14] as translators, and Medmobile paired with Parsinlu [32] [34] models, showed disappointing results, as evidenced by the poor performance displayed in Table VI. Since the translators used in these experiments were not specifically designed for medical translation and were quite small, making them suitable for running on home devices, this approach yielded poor results, nearing random choice in the MMLU and IBMSEE evaluations. Thus, we concluded that this alternative is not viable, and we did not examine any other translators.
Another significant issue with the pipeline alternative is its poor performance in accurately detecting and translating medical terms. This limitation poses a serious challenge, as precise terminology is crucial for effective communication in healthcare settings, where misunderstandings can have serious consequences for patient care and treatment outcomes. The underlying cause of this deficiency is likely due to the fact that the translators employed in these pipeline systems have not been specifically developed for medical translation. Unlike general-purpose translation models, medical translation requires a nuanced understanding of specialized vocabulary, context, and the intricacies of medical language.
Currently, there are no models tailored for medical translation in the Persian language, which means that existing systems are ill-equipped to handle the complexities of medical terminology. As shown in Figure 6, these limitations have resulted in all pipeline alternatives achieving lower win rates against our model, Gaokerena-V. This performance gap highlights the inadequacies of current pipeline approaches in meeting the demands of medical translation, emphasizing the necessity for the development of dedicated medical translator models that can effectively address these specific challenges.
TABLE VI: our model performance in comparison with pipeline alternatives
gao MedMobile MedMobile
kerena-V + gemma2 + parsinlu
@@
VII Ablation study
As an ablation study, we trained qwen2.5 using exactly the same methodology and training setup as employed for aya-expanse. This allowed us to directly compare the performance of the two baseline models under identical conditions, the results of this comparison are presented in Figure 7 and detailed further in Table VII, where it is evident that the trained qwen-2.5 model achieved lower scores across key performance metrics in comparison with gaokerena-V, which utilized Aya-Expanse as its baseline model. These findings highlight the advantages of using aya-expanse as the backbone model and reinforce its effectiveness in producing higher-performing models.
TABLE VII: our model performance in comparison with pipeline alternatives
gaokerena-V Qwen2.5 Qwen2.5
(ours) trained
MMLU-
anatomy(fa) 48.14 41.48 36.29
MMLU-
medical-genetics(fa) 53.0 52.0 50.0
MMLU-
college-medicine(fa) 43.93 43.35 42.77
MMLU-
clinical-knowledge(fa) 55.47 47.92 49.81
MMLU-
professional-medicine(fa) 47.05 43.01 42.27
MMLU-
college-biology(fa) 47.22 42.36 0.50
MMLU(avg) 49.31 45.17 45.35
IBMSEE Sept2023 38.69 33.33 35.71
Number of
parameters 8b 7.6b 7.6b
inference time $\approx 10s$ $\approx 15s$ $\approx 15s$
Figure 7: Our model win rate against trained qwen2.5
@@ VII Ablation study
As mentioned in the baseline model section, we can merge our updated parameters into aya-vision instead of aya-expanse, This will allow us to incorporate medical images, such as MRIs and CT scans, as inputs to our model. However, it is important to note that our models capabilities with multi-modal queries are under-explored. Our future research plans include exploring and improving the model’s responses to queries accompanied by medical images. This will involve fine-tuning the model to better process and align textual and visual data, with the goal of expanding its utility in medical diagnostics and analysis.